Suppose a random variable has population mean -148 and population standard deviation 37.40. What is the lower and upper values of the probability interval containing 95% of the sample means of sample size n = 113?

Lower =  
Upper =  
Answer #1

Ans:

mean=148

population standard deviation=37.40

n=113

lower=148-1.96*(37.40/sqrt(113))=141.10

upper=148+1.96*(37.40/sqrt(113))=154.90